OpenStudy (anonymous):

HELP Please x+y+2=1 2x+4y+2z=8 -x+7y-3z=27

OpenStudy (anonymous):

x = -104/25 y = 79/25 z = -6/25

OpenStudy (anonymous):

\[\color{blue}{x+y+2=1}\] \[\color{red}{x+y=1-2}\] \[\color{brown}{x}\color{orange}+\color{green}y\color{dark green}=\color{black}-\color{black}1\] similarly continue on